A mixture of 2 moles of helium gas (atomic mass = 4 amu) and 1 mole of argon gas (atomic mass = 40 amu) is kept at 300 K in a container. The ratio of the rms speeds

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

v=root(3rt/m)
=>ratio =root(m_argon/m_helium)=root(10)
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty